My math problem says
The student council is sponsoring a carnival to raise money. Tickets cost $5 for adults and $3 for students.The student council wants to raise $450.
a. write an equation to find the number of each type of ticket they should sell.
b. graph your equation
c. use your graph to find two different combinations of tickets sold.
thank you so much and please show your work.

Answer by rwm(914)

5a+3s=450
notice that 3*5=15
15 goes into 450
225 +225
5 into 225=45
3 into 225=75
so we have 5*3* 15+3*5*15=450
5*45+3*75=450
45 adult and 75 student
you graph it and find another